OpenStudy (anonymous):

I have the answer can someone please check me Find (dy/dx) x^2+y^2=8x+8y

OpenStudy (anonymous):

i got dy/dx = (2x-8)/(8-2y)

OpenStudy (anonymous):

2x+2yy'=8+8y' x+yy'=4+4y' (y-4)y'+x=4 (y-4)y'=4-x dy/dx=(4-x)/(y-4)